PHOTOELECTRON SPECTROSCOPY STUDIES OF URANIUM FLUORIDE

\underline{WEI-LI LI}, TIAN JIAN, GARY V. LOPEZ, AND LAI-SHENG WANG , Brown University, Chemistry Department, 324 Brook St, Providence, RI, 02912.

The uranium fluoride anions (UF_{x}^{−}, x=2 - 4) are produced by laser vaporization and investigated using photoelectron spectroscopy at four different photon energies. An extensive vibrational progression of 620 cm^{−1} is observed in the spectra of UF_{4}^{−}, indicating significant geometry change between the anion and the neutral ground states. Franck-Condon simulation is performed to identify the vertical detachment and the 0-0 transition to get the electron affinity of the neutral UF_{4}. Preliminary ab initio calculation shows that the U 5f orbitals participate in the bonding with F 2p orbitals. The UF_{3}^{−} molecule has multiple vibrational modes active upon electron detachment, yielding congested photoelectron spectra. Two vibrational progressions are observed in the UF_{2}^{−} spectra at 580 cm^{−1} and 160 cm^{−1}.